OpenStudy (anonymous):

WILL MEDAL IF ANSWERED The sum of the measures of the interior angles of a convex polygon is 3060 Degrees. How many sides does the polygon have? A. 15 B. 17 C. 18 D. 19

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

how did you get a @Smiley_Funny02

OpenStudy (mathstudent55):

There is a formula for this. \(S = 180(n - 2)\) where S is the sum of the measures of the interior angles of a polygon, and n is the number of sides of the polygon.

OpenStudy (mathstudent55):

For your problem, use the formula above. Replace S with 3060, and solve for n.

OpenStudy (anonymous):

I checked online :)

OpenStudy (mathstudent55):

BTW, @Smiley_Funny02 did two things wrong: 1. He gave an answer without any explanation. 2. His answer is wrong.

OpenStudy (anonymous):

I said will medal if answered

OpenStudy (anonymous):

correctly

OpenStudy (mathstudent55):

@jakematt24 I gave you the formula. Now you can solve it yourself. Just do what I wrote above. Do you still need help?

OpenStudy (anonymous):

Can i just have the answer?

OpenStudy (agl202):

180(n-2) = 3060 n-2 = 3060/180 n-2 = 17 n = 19 sides Answer: D. Hope this helps u!!!

OpenStudy (alexandervonhumboldt2):

now ewe cant give direct ansewers

OpenStudy (anonymous):

Thank You @Agl202

OpenStudy (agl202):

Anytime!!!